OpenStudy (anonymous):

what is 3/5 written as a decimal

OpenStudy (anonymous):

do you have a calculator

OpenStudy (campbell_st):

how many 5s in 100?

OpenStudy (anonymous):

i am doing these questions to give ribbons XD

OpenStudy (anonymous):

still need answer

OpenStudy (aripotta):

once you convert the fraction to have a denominator of 100, you can easily convert it into a decimal

OpenStudy (anonymous):

still need direct answer

OpenStudy (freckles):

answer @campbell_st 's question 100/5=?

OpenStudy (anonymous):

do you know how to do long division? @lordoftanks

OpenStudy (anonymous):

100/5 is 100

OpenStudy (aripotta):

you can't divide a number by another number, and get the same number. unless you divide it by 1. and 5 isn't 1.

OpenStudy (anonymous):

A. 0.3 B. 0.6 C. 0.12 D. 1.67

OpenStudy (freckles):

\[\frac{3}{5} \cdot \frac{\frac{100}{5}}{\frac{100}{5}}=\frac{?}{100}\] if you can figure out 100/5 you are almost there

OpenStudy (freckles):

if you don't know 100/5 do you know at least 30/5?

OpenStudy (anonymous):

a b c or d?

OpenStudy (aripotta):

no one is gonna give you the answer. you have to at least try

OpenStudy (anonymous):

:(